# LoRA Learns Less and Forgets Less

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) is a widely-used parameter-efficient finetuning method for large language models. LoRA saves memory by training only low rank perturbations to selected weight matrices. In this work, we compare the performance of LoRA and full finetuning on two target domains, programming and mathematics. We consider both the instruction finetuning ($\approx$100K prompt-response pairs) and continued pretraining ($\approx$10B unstructured tokens) data regimes. Our results show that, in most settings, LoRA substantially underperforms full finetuning. Nevertheless, LoRA exhibits a desirable form of regularization: it better maintains the base model’s performance on tasks outside the target domain. We show that LoRA provides stronger regularization compared to common techniques such as weight decay and dropout; it also helps maintain more diverse generations. We show that full finetuning learns perturbations with a rank that is 10-100X greater than typical LoRA configurations, possibly explaining some of the reported gaps. We conclude by proposing best practices for finetuning with LoRA.
@@ 1 Introduction
Since its introduction, LoRA has been touted as a strict efficiency improvement that does not compromise accuracy on the new, target domain (Hu et al., 2021; Dettmers et al., 2024; Raschka, 2023; Zhao et al., 2024b). However, only a handful of studies benchmark LoRA against full finetuning for LLMs with billions of parameters, (Ivison et al., 2023; Zhuo et al., 2024; Dettmers et al., 2024), reporting mixed results. Some of these studies rely on older models (e.g. RoBERTa), or coarse evaluation benchmarks (such as GLUE or ROUGE) that are less relevant for contemporary LLMs. In contrast, more sensitive domain-specific evaluations, e.g., code, reveal cases where LoRA is inferior to full finetuning (Ivison et al., 2023; Zhuo et al., 2024). Here we ask: under which conditions does LoRA approximate full finetuning accuracy on challenging target domains, such as code and math?
By training fewer parameters, LoRA is assumed to provide a form of regularization that constrains the finetuned model’s behavior to remain close to that of the base model (Sun et al., 2023; Du et al., 2024). We also ask: does LoRA act as a regularizer that mitigates “forgetting” of the source domain?
In this study, we rigorously compare LoRA and full finetuning for Llama-2 7B (and in some cases, 13B) models across two challenging target domains, code and mathematics. Within each domain, we explore two training regimes. The first is instruction finetuning, the common scenario for LoRA involving question-answer datasets with tens to hundreds of millions of tokens. Here, we use Magicoder-Evol-Instruct-110K (Wei et al., 2023) and MetaMathQA (Yu et al., 2023). The second regime is continued pretraining, a less common application for LoRA which involves training on billions of unlabeled tokens; here we use the StarCoder-Python (Li et al., 2023) and OpenWebMath (Paster et al., 2023) datasets (Table 1).
@@
Figure 1: Learning vs. forgetting tradeoff curves for Llama-2-7B and Llama-2-13B trained on Starcoder-Python. Gray regions are hypothetical Pareto frontiers for performance on the source domain and the code target domain.
We find that for code, LoRA substantially underperforms full finetuning, whereas for math, LoRA closes more of the gap (Sec. 4.1), while requiring longer training. Despite this performance gap, we show that LoRA better maintains source-domain performance compared to full finetuning (Sec. 4.2). Furthermore, we characterize the tradeoff between performance on the target versus source domain (learning versus forgetting). For a given model size and dataset, we find that LoRA and full finetuning form a similar learning-forgetting tradeoff curve: LoRA’s that learn more generally forget as much as full finetuning, though we find cases (for code) where LoRA can learn comparably but forgets less (Sec. 4.3).
We then show that LoRA – even with a less restrictive rank – provides stronger regularization when compared to classic regularization methods such as dropout (Srivastava et al., 2014), and weight decay (Goodfellow et al., 2016). We also show that LoRA provides regularization at the output level: we analyze the generated solutions to HumanEval problems and find that while full finetuning collapses to a limited set of solutions, LoRA maintains a diversity of solutions more similar to the base model (Sun et al., 2023; Du et al., 2024).
Why does LoRA underperform full finetuning? LoRA was originally motivated in part by the hypothesis that finetuning results in low-rank perturbations to the base model’s weight matrix (Li et al., 2018; Aghajanyan et al., 2020; Hu et al., 2021). However, the tasks explored by these works are relatively easy for modern LLMs, and certainly easier than the coding and math domains studied here. Thus, we perform a singular value decomposition to show that full finetuning barely changes the spectrum of the base model’s weight matrices, and yet the difference between the two (i.e. the perturbation) is high rank. The rank of the perturbation grows as training progresses, with ranks 10-100$\times$ higher than typical LoRA configurations (Figure 7).
We conclude by proposing best practices for training models with LoRA. We find that LoRA is especially sensitive to learning rates, and that the performance is affected mostly by the choice of target modules and to a smaller extent by rank.
@@
  • Full finetuning is more accurate and sample-efficient than LoRA in code and math (Sec.4.1).
  • LoRA forgets less of the source domain, providing a form of regularization (Sec. 4.2 and 4.3).
  • LoRA’s regularization is stronger compared to common regularization techniques; it also helps maintaining the diversity of generations (Sec. 4.4).
@@
  • Compared to full finetuning, LoRA is more sensitive to hyperparameters, namely learning rate, target modules, and rank (in decreasing order; Sec. 4.6) .
Dataset Domain # Used Tokens Experiment
StarCoder-Python (Li et al., 2023) Code 20B CPT
OpenWebMath (Paster et al., 2023) Math 8.6B CPT
Magicoder-Evol-Instruct-110K (Wei et al., 2023) Code 72.3M IFT
MetaMathQA (Yu et al., 2023) Math 103M IFT
@@ 2 Background
$\Delta=AB,\quad A\in\mathbb{R}^{d\times r},B\in\mathbb{R}^{r\times k}$
$A_{0}\sim\mathcal{N}(0,1),\quad B_{0}=0.$
The user chooses which $W_{\text{pretrained}}$ to adapt (“target modules”) and the rank $r<<d,k$. By doing so, only $d\times r+r\times k$ parameters are trained instead of $d\times k$, which reduces the memory and FLOPS required for computing the gradient. As an example, applying a $r=16$ LoRA to a 7B weight matrix with $d=k=4096$ trains $<1\%$ of the original parameter count. Sec. D lays out the approximate memory savings by LoRA.
LoRA can additionally enable efficient multi-tenant serving by sharing a single base model across users with personalized adapters (Sheng et al., 2023).
Here, we simplify the combinatorial search over target modules, by grouping modules into three classes: the “Attention” matrices- $\{W_{q}^{(l)},W_{k}^{(l)},W_{v}^{(l)},W_{o}^{(l)}\}_{l=1}^{L}$, the Multi Layer Perceptron (“MLP”) matrices - $\{W_{\text{up}}^{(l)},W_{\text{down}}^{(l)}\}_{l=1}^{L}$, and “All” – their union, for $L$ layers. [^1] The Llama-2 architecture uses the SwiGLU (Shazeer, 2020) architecture for the MLP module of the attention block, which consists of a Gated Linear Unit that uses Swish activation function (Ramachandran et al., 2017)). SwiGLU has a gating matrix $W_{\text{gate}}$ in addition to $W_{\text{up}}$ and $W_{\text{down}}$, which we did not modify in our LoRA experiments. Note that while in the original study LoRA was only applied to the attention matrices (Hu et al., 2021), it has become the standard to target additional modules (Raschka, 2023; Dettmers et al., 2024).
@@ 3 Experimental Setup
We train on code and math datasets that have been shown to increase downstream performance. We motivate the training datasets and evaluation benchmarks below.
@@ 3.1 Datasets for Continued Pretraining (CPT) and Instruction Finetuning (IFT)
This dataset (Li et al., 2023) consists of permissively licensed repositories from GitHub, including Git commits, in 80+ programming languages. We chose the Python subset and sub-sampled it to 20B tokens.
@@
We trained on a subset of up to 8.59B out of 14.7B tokens. The dataset (Paster et al., 2023) includes mathematical web pages from Common Crawl, correctly formatted to preserve mathematical content such as LaTeX equations. [^2] https://huggingface.co/datasets/open-web-math/open-web-math We note that this dataset contains a considerable amount of full English sentences.  [^3] Out of a random selection of 100K examples, a regex search shows that 75% of the examples contain LaTex. The data is classified as 99.7% English and ”overwhelmingly English” by the langdetect and fasttext tools.
@@
This dataset (Wei et al., 2023) contains 72.97M tokens of programming questions and answers. It reproduces the “Evol-Instruct’ dataset of WizardCoder (Luo et al., 2023): an LLM (GPT-4) is iteratively prompted to increase the difficulty of a set of question-answer pairs (from Code Alpaca; Chaudhary (2023)).
@@
This dataset (Yu et al., 2023) was built by bootstrapping mathematical word problems from the training sets of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) by rewriting the questions with variations using GPT-3.5. This dataset contains 395K question-answer pairs and roughly 103M tokens. [^4] https://huggingface.co/datasets/meta-math/MetaMathQA
@@ 3.2 Measuring Learning with Coding and Math Benchmarks (target domain evaluation)
This benchmark (Chen et al., 2021)) contains 164 problems that involve generating of a Python program given a docstring and a function signature. A generation is considered correct if it passes all supplied unit tests. We use Code Generation LM Evaluation Harness (Ben Allal et al., 2022), configured to output 50 generations per problem, sampling with softmax temperature=0.2, and calculating “pass@1”.
@@
This benchmark (Cobbe et al., 2021) includes a collection of 8.5K grade-school math word problems . We evaluate on the test split of GSM8K (1,319 samples) as implemented in LM Evaluation Harness (Gao et al., 2023), with default generation parameters (temperature=0, five few-shot, pass@1).
@@ 3.3 Forgetting Metrics (source domain evaluation)
HellaSwag
This benchmark (Zellers et al., 2019) includes 70K problems, each describing an event with multiple possible continuations. The task is to pick the most plausible continuation, which requires making inferences about nuanced everyday situations.
WinoGrande
This benchmark (Sakaguchi et al., 2019) also assesses commonsense reasoning. It includes 44K problems with sentences that require ambiguous pronoun resolution.
ARC-Challenge
This benchmark (Clark et al., 2018) consists of 7,787 grade-school level, multiple-choice science questions, testing capabilities in complex reasoning and understanding of scientific concepts.
@@ 4 Results
4.1 LoRA underperforms full finetuning in programming and math tasks
@@
We compare LoRA and full finetuning after performing an exhaustive learning rate sweep for each method, which we found to be crucial (Dettmers et al., 2024). We include learning rate sweep results in Figure 8.
@@
We perform a sample-efficiency analysis – i.e., compute the learning metrics as a function of training samples seen – for both LoRA and full finetuning. For IFT, we train separate models for $1,2,4,8,16$ epochs. For CPT, we manipulate the number of unique tokens ($0.25,0.5,1,2,4,8,16,20$ billion), using individual learning rate cooldown schedules. We train six LoRA models for each condition (3 target modules [“Attention”, “MLP”, and “All”] $\times$ 2 ranks [$16,256$]).
@@
Figure 2: LoRA underperforms full finetuning in programming and in math. We use six LoRA configurations per condition and plot their mean, minimum, and maximum values in purple. (A) Starcoder-Python, (B) Magicoder-Evol-Instruct-110K, (C) OpenWebMath, (D) MetaMathQA. Note that 16 epochs are $\approx$1.16B and $\approx$1.6B tokens, for Magicoder-Evol-Instruct-110K and MetaMathQA.
In Fig. 2, we summarize the performance of the six LoRA models with their minimum, average, and maximum performance for each training duration (in purple), and compare them to full finetuning (solid black lines) and the base model (dashed horizontal line); the results are further broken down per LoRA configuration in Fig. S2. We first note that for both domains, IFT leads to more significant improvements compared to CPT, which is expected because IFT’s problems are more similar to the evaluation problems (e.g., for code, IFT achieves maximum HumanEval of 0.50 VS 0.26 for CPT).
@@
For Code CPT (Fig. 2A), we identify a substantial gap between full finetuning and LoRA that grows with more data. The overall best LoRA peaks at 16B tokens (rank=256, “All”) with HumanEval=0.175, roughly matching full finetuning with 1B tokens (HumanEval=0.172). Full finetuning reaches its peak HumanEval of 0.263 at 20B tokens. For Code IFT (Fig. 2B), the best performing LoRA ($r=256$, “All”) achieves HumanEval=0.407 at epoch 4, meaningfully underperforming full finetuning at epoch 2 (HumanEval=0.464) and at its peak HumanEval score of 0.497 at 8 epochs. Math CPT (Fig. 2C) results show that training for 1B tokens or fewer degrades GSM8K results below baseline (GSM8K=0.145). Improvements appear with more data, where the best LoRA (rank=256, “All”) achieves GSM8K=0.187 at 8.6 billion tokens, underperforming full finetuning at 4.3 billion tokens (GSM8K=0.191) and at 8.6 billion tokens (GSM8K=0.230). LoRA closes most of the gap with full finetuning in the Math IFT (Fig. 2D) dataset. However, LoRA still remains less sample efficient. LoRA ($r=256$, “All”) peaks at 4 epochs (GSM8K=0.622) while full finetuning achieves GSM8K=0.640 at 2 epochs and peaks at 4 epochs, with GSM8K=0.642.666We note that the original MetaMath paper reports a maximum accuracy of 0.665 when (fully) finetuning Llama-2-7B on the MetaMathQA dataset. We attribute this to small differences in hyperparameters; they trained on 3 epochs with a batch size of 128 using the AdamW optimizer, a learning rate of 2e-5, a learning rate warmup of 3%. Both methods substantially exceed the base model. We hypothesize that the milder gaps here correspond to a smaller domain shift between the math problems and the pretraining data, different from the larger shifts in code.
In summary, across LoRA configurations and training durations, it still appears to underperform full finetuning. These effects are more pronounced for programming than math. For both domains, instruction finetuning leads to larger accuracy gains than continued pretraining.

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) is a widely-used parameter-efficient finetuning method for large language models. LoRA saves memory by training only low rank perturbations to selected weight matrices. In this work, we compare the performance of LoRA and full finetuning on two target domains, programming and mathematics. We consider both the instruction finetuning ($\approx$100K prompt-response pairs) and continued pretraining ($\approx$20B unstructured tokens) data regimes. Our results show that, in the standard low-rank settings, LoRA substantially underperforms full finetuning. Nevertheless, LoRA better maintains the base model’s performance on tasks outside the target domain. We show that LoRA mitigates forgetting more than common regularization techniques such as weight decay and dropout; it also helps maintain more diverse generations. Finally, we show that full finetuning learns perturbations with a rank that is 10-100$\times$ greater than typical LoRA configurations, possibly explaining some of the reported gaps. We conclude by proposing best practices for finetuning with LoRA.
@@ 1 Introduction
LoRA is widely adopted for finetuning LLMs under hardware constraints, but the jury is still out on whether it compromises performance compared to full finetuning. The two seminal methods papers on the topic, which introduce LoRA (Hu et al., 2021) and its more recent combination with model quantization (QLoRA; Dettmers et al. (2024)), reported that LoRA performs better or equivalent to full finetuning. More empirical work (Ghosh et al., 2024; Zhao et al., 2024b) reaches a similar conclusion; this sentiment is echoed in an array of industry blog posts as well (e.g., Raschka (2023); Niederfahrenhorst et al. (2023)). At the same time, there is evidence that LoRA underperforms full finetuning (Ivison et al., 2023; Zhuo et al., 2024), and the need to improve upon LoRA has led to the development of enhanced LoRA variants (Hayou et al., 2024; Meng et al., 2024; Li et al., 2023b; Shi et al., 2024) or alternative low-rank approximation methods (e.g Liu et al. (2024); Zhao et al. (2024a)). To shed light on this ongoing debate, we ask: under which conditions does LoRA approximate full finetuning accuracy on challenging target domains, such as code and math?
By training fewer parameters, LoRA is hypothesized to constrain the finetuned model from diverging significantly from the base model (Sun et al., 2023; Du et al., 2024). This potential characteristic is particularly helpful for LLM finetuning, a form of continual learning where specializing in new domains can come at the expense of base model capabilities (Wang et al., 2024) (a phenomenon known its extreme form as “catastrophic forgetting” McCloskey & Cohen (1989); French (1999)). To date, only a few studies have examined forgetting in modern LLMs (Vu et al., 2022; Kleiman et al., 2023; Kalajdzievski, 2024). To address this gap, we also ask: when performing continual learning on a new domain, to what extent does LoRA mitigate forgetting of base model capabilities?
In this study, we compare LoRA and full finetuning for Llama-2-7B models across two challenging target domains, code and mathematics. Within each domain, we explore two training regimes. The first regime is continued pretraining, which involves training on billions of unlabeled domain-specific tokens, most commonly via full finetuning; here we use the StarCoder-Python (Li et al., 2023a) and OpenWebMath (Paster et al., 2023) datasets (Table 1). The second is instruction finetuning, the common scenario for LoRA involving question-answer datasets with tens to hundreds of millions of tokens. Here, we use Magicoder-Evol-Instruct-110K (Wei et al., 2023) and MetaMathQA (Yu et al., 2023).
@@
We find that with commonly used low-rank settings, LoRA substantially underperforms full finetuning, while typically requiring longer training (Sec. 4.1). In continued pretraining, the performance gap between full finetuning and LoRA is not closed even with high ranks. In instruction finetuning, on the other hand, high ranks can match full finetuning performance.
Despite LoRA’s limitations, we show that it consistently maintains better source-domain performance compared to full finetuning (Sec. 4.2). Furthermore, we characterize the tradeoff between learning and forgetting (Sec. 4.3). We then show that LoRA – even with higher rank – mitigates forgetting more aggressively than classic regularization techniques that aim to prevent overfitting, such as dropout (Srivastava et al., 2014; Goodfellow et al., 2013), and weight decay (Goodfellow et al., 2016). Moreover, by analyzing the generated solutions to HumanEval problems, we demonstrate that while full finetuning tends to produce a limited set of solutions, LoRA produces a wider range of solutions more akin to those of the base model (Sun et al., 2023; Du et al., 2024)
Why does LoRA underperform full finetuning? LoRA was originally motivated in part by the hypothesis that finetuning results in low-rank perturbations to the base model’s weight matrix (Li et al., 2018; Aghajanyan et al., 2020; Hu et al., 2021). However, the tasks explored by these prior works are relatively easy for modern LLMs, and certainly easier than the coding and math domains studied here. Thus, we perform a singular value decomposition to show that full finetuning barely changes the spectrum of the base model’s weight matrices, and yet the difference between the two (i.e. the perturbation) is high rank. The rank of the perturbation grows as training progresses, with ranks 10-100$\times$ higher than typical LoRA configurations (Figure 6).
We conclude by proposing best practices for training models with LoRA. We find that LoRA is very sensitive to hyperparameters, including learning rates, choice of target modules, ranks, and scaling factors; setting these properly is a prerequisite to approach full finetuning performance.
@@
  • Full finetuning is more accurate and sample-efficient than LoRA in continued pretraining (CPT) for code and math; in instruction finetuning (IFT), higher ranks can close most of the gaps (Sec.4.1).
  • LoRA forgets less of the source domain (Sec. 4.2 and 4.3).
  • LoRA forgets less than common regularization techniques; it also helps maintaining the diversity of generations (Sec. 4.5).
@@
  • A hyperparameter sensitivity analysis for LoRA, as well as practical recommendations (Sec. 4.7).
Model checkpoints and LoRA adapters can be accessed at https://github.com/danbider/lora-tradeoffs.
Code Math
CPT StarCoder-Python (up to 20B tokens) OpenWebMath (14.7B tokens)
IFT Magicoder-Evol-Instruct-110K (72.97M tokens) MetaMathQA (103M tokens)
@@ 2 Background
$\Delta=\gamma_{r}AB,\quad A\in\mathbb{R}^{d\times r},\quad B\in\mathbb{R}^{r\times k}.$
Most common implementations initialize $A_{0}\sim\mathcal{N}(0,1),~{}B_{0}=0$ and set the scalar $\gamma_{r}=\alpha/r$ with a controllable hyperparameter $\alpha$. The user chooses which $W_{\text{pretrained}}$ to adapt (“target modules”), the rank $r<<d,k$, and the hyperparameter $\alpha$. By doing so, only $d\times r+r\times k$ parameters are trained per module instead of $d\times k$, which reduces the memory and FLOPS required for computing the gradient. As an example, applying a $r=16$ LoRA adapter to a 7B weight matrix with $d=k=4096$ trains $<1\%$ of the original parameter count. Appendix Sec. H lays out the approximate memory savings by LoRA during training.
LoRA’s introduction and first applications targeted only the $W_{q}$ and $W_{v}$ matrices in the self-attention module (Hu et al., 2021). Since then, it has become best practice to target all transformer modules (Raschka, 2023; Dettmers et al., 2024), i.e., $\{W_{q}^{(l)},W_{k}^{(l)},W_{v}^{(l)},W_{o}^{(l)}\}_{l=1}^{L}$ in the self-attention modules, and $\{W_{\text{gate}}^{(l)},W_{\text{up}}^{(l)},W_{\text{down}}^{(l)}\}_{l=1}^{L}$ in the feedforward modules for $L$ layers in, say, a Llama architecture (Hu et al., 2021; Touvron et al., 2023).
@@ 3 Experimental Setup
We train on code and math datasets that have been shown to increase downstream performance. We motivate the training datasets and evaluation benchmarks below. All training was done using the Databricks MosaicML composer [^1] https://github.com/mosaicml/composer, streaming [^2] https://github.com/mosaicml/streaming, and llm-foundry [^3] https://github.com/mosaicml/llm-foundry repositories, as well as the HuggingFace peft library.
@@ 3.1 Datasets for Continued Pretraining (CPT) and Instruction Finetuning (IFT)
(Li et al., 2023a) This dataset consists of permissively licensed repositories from GitHub, including Git commits, in 80+ programming languages. We chose the Python subset and sub-sampled it to 20B tokens.
@@
(Paster et al., 2023) This dataset contains 14.7B tokens derived from mathematical web pages from Common Crawl, correctly formatted to preserve mathematical content such as LaTeX equations. [^4] https://huggingface.co/datasets/open-web-math/open-web-math To match with the StarCoder-Python dataset, we trained on up to 20B tokens, repeating tokens beyond the first 14.7B. An analysis of this dataset shows that it contains a considerable amount of full English sentences. [^5] Out of a random selection of 100K examples, a regex search shows that 75% of the examples contain LaTex. The data is classified as 99.7% English and “overwhelmingly English” by the langdetect and fasttext tools.
@@
(Wei et al., 2023) This dataset contains 72.97M tokens of programming questions and answers. It reproduces the “Evol-Instruct” dataset of WizardCoder (Luo et al., 2023b) by iteratively prompting an LLM (GPT-4) to increase the difficulty of a set of question-answer pairs from Code Alpaca (Chaudhary, 2023). [^6] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K
@@
(Yu et al., 2023) This dataset was built by bootstrapping mathematical word problems from the training sets of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) by rewriting the questions with variations using GPT-3.5. This dataset contains 395K question-answer pairs and roughly 103M tokens. [^7] https://huggingface.co/datasets/meta-math/MetaMathQA
@@ 3.2 Measuring Learning with Coding and Math Benchmarks (target domain evaluation)
(Chen et al., 2021) This benchmark contains 164 problems that involve generating a Python program given a docstring and a function signature. A generation is considered correct if it passes all supplied unit tests. We use the Code Generation LM Evaluation Harness (Ben Allal et al., 2022) configured to output 50 generations per problem, and calculate “pass@1” with softmax temperature=0.2 and top_p=0.95 for 0-shot HumanEval.
@@
(Cobbe et al., 2021) This benchmark includes a collection of 8.5K grade-school math word problems. We evaluate on the test split of GSM8K (1,319 samples) as implemented in the LM Evaluation Harness (Gao et al., 2023), with default generation parameters (temperature=0, 5 few-shot, pass@1).
@@ 3.3 Forgetting Metrics (source domain evaluation)
We use the following benchmarks to asses degradation of base model capabilities. HellaSwag (Zellers et al., 2019) includes 70K problems that describe an event with multiple possible continuations. The task is to pick the most plausible continuation, which requires making inferences about nuanced everyday situations. WinoGrande (Sakaguchi et al., 2019) also assesses commonsense reasoning. It includes 44K problems with sentences that require ambiguous pronoun resolution. ARC-Challenge (Clark et al., 2018) consists of 7,787 grade-school level, multiple-choice science questions, and tests complex reasoning and understanding of scientific concepts. These benchmarks involve multiple-choice questions that use the predicted logits for calculating accuracy, and do not require specifying further generation hyperparameters. All forgetting metrics were computed using the MosaicML Gauntlet evaluation harness (Dohmann, 2023). [^9] https://github.com/mosaicml/llm-foundry/tree/main/scripts/eval
@@ 4 Results
4.1 Target-domain performance: LoRA at low ranks underperforms full finetuning
@@
We compare LoRA and full finetuning after performing an exhaustive learning rate sweep for each method, which we found to be crucial (Dettmers et al., 2024). We include learning rate sweep results in Figure S1.
@@
We perform a sample-efficiency analysis – i.e., compute the learning metrics as a function of training samples seen – for both LoRA and full finetuning. For IFT, we train separate models for $1,2,4,8$, and $16$ epochs. For CPT, we vary the number of training tokens ($0.25,0.5,1,2,4,8,16,20$ billion), using individual learning rate cooldown schedules. For each condition, we train one full finetuning model and three LoRA models with ranks $r=16,64,256$ noting that most LoRA papers use a “low” rank of 8-64, (e.g., Dettmers et al. (2024); Zhuo et al. (2024)). The LoRA models target all transformer modules and use $\alpha=2r$, as known to be best practice (Raschka, 2023). For further details on experimental setup and hyperparameters, see Appendix Sec. A.
@@
Figure 1: LoRA performance scales by rank and underperforms full finetuning in code and math. (A) Starcoder-Python, (B) Magicoder-Evol-Instruct-110K, (C) OpenWebMath, (D) MetaMathQA. In (A) and (B) $y$-axis: HumanEval pass@1. In (C) and (D) $y$-axis: GSM8K strict match. In all panels, “base model” indicates Llama-2-7B without instruction finetuning. Note that 16 epochs are $\approx$1.16B and $\approx$1.6B tokens, for Magicoder-Evol-Instruct-110K and MetaMathQA, respectively.
The results appear in Fig. 1. We first note that for both programming and math, IFT improves evaluation scores much more than CPT, which is expected because the samples in each IFT dataset are more similar to the evaluation problems (e.g., for code, IFT achieves maximum HumanEval of 0.497 vs. 0.263 for CPT).
For Code CPT (Fig. 1A and Table S1), we identify a substantial gap between full finetuning and LoRA that grows with more data. The best LoRA model, with rank $r=256$, peaks at 20B tokens with HumanEval=0.224, roughly matching full finetuning with 4B tokens (HumanEval=0.218). Full finetuning reaches its peak HumanEval of 0.263 at 20B tokens. A clear ordering by rank emerges after the initial 1B CPT tokens.
For Code IFT (Fig. 1B and Table S5), HumanEval accuracy is clearly ordered by rank from the very first epoch. The more common $r=16$ and $r=64$ LoRA configurations have lower accuracy than full finetuning, with HumanEval scores of 0.358 and 0.417 at epoch 4, respectively). With a high LoRA rank ($r=256$), full finetuning performance can be matched (LoRA=0.498 in epoch 4, full finetuning=0.497 in epoch 8). In Appendix Sec. F we perform a more sensitive HumanEval analysis, calculating pass@$k$ as a function of $k=1,\dots,256$ with a higher temperature of 0.8 for full finetuning and the LoRA models (at epoch 4). This analysis shows that full finetuning is superior to $r=256$ for $k<64$, after which the two are equal.
Math CPT (Fig. 1C and S3) results closely echo those of code CPT. Consistent patterns in GSM8K emerge at 4B tokens. Full finetuning opens a gap in GSM8K which widens with more data. Similarly, LoRA performance is ordered by rank. The best LoRA ($r=256$) peaks at 16B tokens (GSM8K=0.203), underperforming full finetuning at 4B tokens (GSM8K=0.224) and at its peak at 20B tokens (GSM8K=0.293).
@@
LoRA closes much of the gap with full finetuning in the Math IFT (Fig. 1D and Table S7) dataset, while remaining less sample efficient. Both methods substantially improve upon the base model; LoRA ($r=256$) peaks at 8 epochs (GSM8K=0.634) while full finetuning achieves GSM8K=0.641 at 2 epochs and peaks at 4 epochs, with GSM8K=0.642.101010We note that the original MetaMath paper reports a maximum accuracy of 0.665 when (fully) finetuning Llama-2-7B on the MetaMathQA dataset. We attribute this to small differences in hyperparameters; they trained on 3 epochs with a batch size of 128 using the AdamW optimizer, a learning rate of 2e-5, a learning rate warmup of 3%. Unlike the code IFT dataset, $r=64$ suffices to approach full finetuning and achieve GSM8K=0.624 at epoch 4. We suggest that lower ranks are effective here because English mathematics problems involve a smaller domain shift from the pretraining data as compared to coding ones.
In summary, in CPT, LoRA underperforms full finetuning across all configurations. In IFT, and especially in code, high LoRA ranks are required to close the gap with full finetuning.